# Spherically symmetric solutions to the Einstein-scalar field conformal constraint equations

### v1

@@ Appendix A Explicit solutions on the sphere
$N^{2}(\varphi^{\prime})^{2}-n^{2}\varphi^{2}+2nN\cot(r)\varphi\varphi^{\prime}+{1\over\sin^{n}(r)}\int_{0}^{r}{I(V,\psi,\rho,\varphi)\over\varphi^{2N}}(\varphi^{N}\sin^{n}(s))^{\prime}\,ds>0.$
@@
${1\over\sin^{n}(r)}\int_{0}^{r}{I(V,\psi,\rho,\varphi)\over\varphi^{2N}}(\varphi^{N}\sin^{n}(s))^{\prime}\,ds=\phi,$ (A.1)
@@
$(\varphi^{N}\sin^{n}(r))^{\prime}=\bigg({n-1\over n}\bigg){(\phi\sin^{n}(r))^{\prime}(\varphi^{N}\sin^{n}(r))^{2}\over\rho^{2}\sin^{2n}(r)}.$
@@
$\varphi^{N}=-\bigg({n\over n-1}\bigg){1\over\sin^{n}(r)\big(\int_{\pi/2}^{r}{(\phi\sin^{n}(s))^{\prime}\over\rho^{2}\sin^{2n}(s)}\,ds+c_{1}\big)},$
@@
$\varphi^{N}=\bigg({n\over n-1}\bigg){\rho\over 1-c\rho\sin^{n}(r)},$
@@
$\varphi^{\prime}={1\over N}\bigg({n\over n-1}\bigg)^{1/N}{\rho^{\prime}+nc\rho^{2}\cos(r)\sin^{n-1}(r)\over\rho^{1-1/N}\big(1-c\rho\sin^{n}(r)\big)^{1+1/N}}.$
@@
it follows that $\varphi\sim\bigg({n\over n-1}\bigg)^{1/N}\rho^{1/N}$ and $\cot(r)\varphi^{\prime}\varphi$ is either positive or small negative. Hence, by definition, we get from (A.1A.2) that

### v2

@@ Appendix A Explicit solutions on the sphere
$N^{2}(\varphi^{\prime})^{2}-n^{2}\varphi^{2}+2nN\cot(r)\varphi\varphi^{\prime}+\frac{1}{\sin^{n}(r)}\int_{0}^{r}\frac{I(V,\psi,\rho,\varphi)}{\varphi^{2N}}(\varphi^{N}\sin^{n}(s))^{\prime}\,ds>0.$
@@
$\frac{1}{\sin^{n}(r)}\int_{0}^{r}\frac{I(V,\psi,\rho,\varphi)}{\varphi^{2N}}(\varphi^{N}\sin^{n}(s))^{\prime}\,ds=\phi,$ (A.1)
@@
$(\varphi^{N}\sin^{n}(r))^{\prime}=\big(\tfrac{n-1}{n}\big)\frac{(\phi\sin^{n}(r))^{\prime}(\varphi^{N}\sin^{n}(r))^{2}}{\rho^{2}\sin^{2n}(r)}.$
@@
$\varphi^{N}=-\big(\tfrac{n-1}{n}\big)\frac{1}{\sin^{n}(r)\big(\int_{\pi/2}^{r}\frac{(\phi\sin^{n}(s))^{\prime}}{\rho^{2}\sin^{2n}(s)}\,ds+c_{1}\big)},$
@@
$\varphi^{N}=\big(\tfrac{n}{n-1}\big)\frac{\rho}{1-c\rho\sin^{n}(r)},$
@@
$\varphi^{\prime}=\frac{1}{N}\bigg(\frac{n}{n-1}\bigg)^{1/N}\frac{\rho^{\prime}+nc\rho^{2}\cos(r)\sin^{n-1}(r)}{\rho^{1-1/N}\big(1-c\rho\sin^{n}(r)\big)^{1+1/N}}.$
@@
it follows that $\varphi\sim\big(\tfrac{n}{n-1}\big)^{1/N}\rho^{1/N}$ and $\cot(r)\varphi^{\prime}\varphi$ is either positive or negative and small. Hence, by definition, we get from (A.1) and (A.2) that